OpenStudy (anonymous):

A batch of 500 containers for frozen orange juice contain 5 that are defective. 2 are selected at random without replacement. A=event that 1st is defective, B=event that 2nd is defective, C=event that 3rd is defective. a) What is the probability that the 2nd one selected is defective given that 1st one was defective? b) What is the probability that both are defective? c) What is probability that both are acceptable? d) Three containers are selected without replacement. What is the probability that the 3rd one selected is defective given 1st & 2nd are defective?

OpenStudy (anonymous):

and these two more parts as well e) What is the probability that 3rd is defective given 2nd defective & 1st acceptable? f) What is the probability that all 3 are defective?

OpenStudy (anonymous):

n = number of trials r = number of success (n Choose r) p^r (1 - p)^(n-r)

OpenStudy (anonymous):

a) 4 defectives left in 499 => 4/499 (b) 5/500 *4/499 = 1/12475 (c) 495/500 *494/499 = ...... (d) 3 defectives are left out of 498 => 3/498 (e) 4 defectives are left in 498 => 4/498 = ..... (f ) 5/500 *4/499 *3/498= ......